Your goal is to retire with $500,000. If you have realized $75,250 from selling off your
tech portfolio and have invested it in government bonds that yield 4.75% annually, what
is the closest number of years until you can retire?

Answer by mananth(16949)

Principal P = 75250
Amount= 500000
n= number of periods
compounded (t) 1
Rate = 4.25 0
n = ln(FV / PV) / ln(1 + r) (r/q))

n=( ln ( 500000 / 75250 )/ln( 1 + 0.04 )
n=( 1.89 )/ ln 1.04
n=( 1.89 / 0.04 )
n= 45.5
46 years